OpenStudy (anonymous):

nine is 3 times a number and two.

OpenStudy (anonymous):

nine is three times a the difference between a number and two*

OpenStudy (anonymous):

is that 3+ 2? =5? since 3 is a square of 9?

OpenStudy (anonymous):

9= 3n+ 2 If by "and" you mean plus.

OpenStudy (anonymous):

9= 3(x -2) 9= 3x -6 +6 +6 15=3 15/3 = 5 x=5